OpenStudy (sissyedgar):

Help please, Medals! :) What is the reciprocal of 1/5? A. -5 B. -1/5 C. 1/5 D. 5

OpenStudy (sissyedgar):

@Bookworm14

OpenStudy (kl0723):

D=5

OpenStudy (sissyedgar):

ok, thx

OpenStudy (kl0723):

just flip it

OpenStudy (sissyedgar):

Express your answer in lowest terms. Divide 1/5 {divided by} 1/3 = ?

OpenStudy (sissyedgar):

@Bookworm14

OpenStudy (bookworm14):

3/5 :)

OpenStudy (sissyedgar):

ty :), i have a few more...

OpenStudy (kl0723):

\[\frac{ 1 }{ 5 } \div \frac{ 1 }{ 3 } = \frac{ 1 }{ 5 } \times \frac{ 3 }{ 1 }\]

OpenStudy (bookworm14):

ok :)

OpenStudy (bookworm14):

Yikes! I gotta go but I will be back on later, if you need me just tag me in a post or msg me :) ~Bree

OpenStudy (sissyedgar):

^^^ what is 15 {divided by} 3/5?

OpenStudy (sissyedgar):

A. -25 B. -9 C. 1/9 D. -1/25

OpenStudy (sissyedgar):

wait a sec, imma open this question in a diffrent tab thingie! :P